OpenStudy (anonymous):

10–9 = 1.000000001 0.0000000001 0.000000001 0.0000001

OpenStudy (anonymous):

None of the above?

OpenStudy (anonymous):

10-9=1

geerky42 (geerky42):

I'm sure he means \(\Large 10^{-9}\)

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

C then.

OpenStudy (anonymous):

You just go back 9 places.

OpenStudy (anonymous):

thnk you .

OpenStudy (anonymous):

10^1=10 10^0=1 10^(-1)=0.1 10^(-2)=0.01 10^(-3)=0.001 See the pattern?

OpenStudy (anonymous):

Your welcome.

geerky42 (geerky42):

\(10^0=1\) ~ no zero \(10^{-1}=0.1\) ~ one zero \(10^{-2}=0.01\) ~ two zeros \(10^{-3}=0.001\) ~ three zeros \(10^{-4}=0.0001\) ~ four zeros \(\vdots\) \(\vdots\) \(\vdots\)